how many prime factors are their of number 666666 ?
Answer by Edwin McCravy(20054)

666666÷2 = 333333

333333÷3 = 111111

111111÷3 = 37037

37037÷7 = 5291

5291÷11 = 401

481÷13 = 37

Therefore 666666 = 2*3*3*7*11*13*37

Prime factors are 2,3,7,11,13,37

Answer: 666666 has 6 prime factors.

Edwin